30 years hence a man will be twice as old as he will be 3 years hence. How old is he now?
Answer by ankor@dixie-net.com(22740)

Let n = man's present age
:
Just write an equation for what it says:
"30 years hence a man will be twice as old as he will be 3 years hence."
n + 30 = 2(n + 3)
n + 30 = 2n + 6
30 - 6 = 2n - n
n = 24 yrs is his present age
:
:
Check this in the original statement:
24 + 30 = 2(24 + 3)
54 = 2(27)